OpenStudy (anonymous):

How would I find the coterminal angle of -960 that is between 0 and 360 degrees? And also coterminal for 29pi/12 that is between 0 and 2pi.

OpenStudy (ranga):

Add multiples of 360 degrees until you get the angle between 0 and 360 degrees. Adding or subtracting multiples of 360 degrees (full rotation) will end up in the same place.

OpenStudy (anonymous):

so I would multiply 360 by multiples of itself until i get to positive 960?

OpenStudy (ranga):

-960 + 360 = -600 -600 + 360 = -240 -240 + 360 = 120. 120 degrees, -240 degrees, -600 degrees, -960 degrees are all coterminal.

OpenStudy (anonymous):

ohh okay bingo! Thanks

OpenStudy (ranga):

You are welcome.